40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 15 15:10:48 2015, gautam, ericq, rana, Summary, LSC, PRFPMI lock & various to-do's powerSpec0915.pdf
    Reply  Tue Sep 15 18:35:21 2015, ericq, Summary, LSC, some further notes 
       Reply  Tue Sep 15 20:44:13 2015, gautam, Summary, LSC, Checking the delay line phase shifter DS050339 IMG_20150915_193100.jpgCalibration.pdf
       Reply  Thu Sep 17 03:48:10 2015, ericq, Summary, LSC, some further notes 
          Reply  Thu Sep 17 13:06:05 2015, ericq, Summary, LSC, Low input impedance on CM board 
Message ID: 11599     Entry time: Tue Sep 15 15:10:48 2015     Reply to this: 11601
Author: gautam, ericq, rana 
Type: Summary 
Category: LSC 
Subject: PRFPMI lock & various to-do's 
I was observing Eric while he was attempting to lock the PRFPMI last night. The handoff from ALS to LSC was not very smooth, and Rana suggested looking at some control signals while parked close to the PRFPMI resonance to get an idea of what frequency bands the noise dominated in. The attached power spectrum was taken while CARM and DARM were under ALS control, and the PRMI was locked using REFL_165. The arm power was fluctuating between 15 and 50. Most of the power seems to be in the 1-5Hz band and the 10-30Hz band.

Rana made a number of suggestions, which I'm listing here. Some of these may directly help the above situation, while the others are with regards to the general state of affairs.

  • Reroute both (MC and arm) FF signals to the SUS model
  • For MC, bypass LSC
  • Rethink the MC FF -
  • Leave the arm FF on all the time?
  • The positioning of the accelerometer used for MC FF has to be bettered - it should be directly below the tank
  • The IOO model is over-clocking - needs to be re-examined
  • Fix up the DC F2P - Rana mentioned an old (~10 yr) script called F2P ratio, we should look to integrate the Python scripts used for lock-in/demod at the sites with this
  • Look to calibrate MC_F
  • Implement a high BW CARM servo using ALS
  • Gray code implementation for EPICS gain-stepping

Attachment 1: powerSpec0915.pdf  75 kB  | Hide | Hide all
powerSpec0915.pdf
ELOG V3.1.3-